Determine the average function for each of the following total functiona.

a) Total revenue= 100Q-Q2

b) Total Cost= 1000+ 10Q+0.01Q2

c) Total Profit = 50Q - 0.1Q2 - 1000


1
Expert's answer
2021-07-14T11:59:33-0400

a) AVR=TR\Q

=(100Q-Q²)\Q

=100+Q

b)AVC=TC\Q

=(1000+10Q+0.01Q²)\Q

=1000\Q+10+0.01Q

c)AVP=TP\Q

=(50Q-0.1Q²-1000)\Q

=50-0.1Q-1000/Q